OpenStudy (anonymous):

Can someone please help me graph this rational function: 2 ---------- x^2-2x-3 Also need the Domain: x and y intercepts: horizontal asymptotes: vertical asymptotes: *This is a practice problem, if someone can help me find the answers to this problem i can help myself do the actual problem that i need. Thanks!

hero (hero):

\[y = \frac{2}{x^2 - 2x - 3}\]

hero (hero):

First factor the denominator \[y = \frac{2}{(x - 3)(x + 1)}\]

OpenStudy (anonymous):

Okay

hero (hero):

Next, figure out the asymptotes. We know \(x \ne 3\) and \(x \ne 1\)

OpenStudy (anonymous):

So the vertical asymptotes would be x=3 , x=-1?

hero (hero):

The way to know if those are asymptotes is to test numbers very close such x = 2.9 and x = 3.1 if they diverge, meaning if one is negative and the other is positive, then it is safe to assume that x = 3 is an asymptote. Same goes for x = -1

hero (hero):

If you know how to program your calculator, you can program the function into the calculator to test such points.

OpenStudy (anonymous):

Okay, so vertical asymptotes are x=3 and x=-1. Now how about the rest?

hero (hero):

Okay, yes, they are both asymptotes, so we can plot those.

hero (hero):

To find the horizontal asymptotes, we have to figure out what number we can input into y such that we cannot solve for x.

hero (hero):

There's one obvious number that would work.

OpenStudy (anonymous):

0?

hero (hero):

Yes, very good. We at least know that it is not part of the graph.

hero (hero):

There's no other number that will work

OpenStudy (anonymous):

Okay, so the horizontal asymptote is y=0?

hero (hero):

Let's try to find some intercepts.

hero (hero):

What happens if x = 0?

OpenStudy (anonymous):

Ive already tried, are these correct. x intercept, x= (0, -2/5) y intercept, (2,0) ?

hero (hero):

For the y intercept, you insert x = 0, then solve for y

hero (hero):

For the x intercept, you insert y = 0, then try to solve for x. We already tried this and we didn't get anywhere, so we can assume there is no x intercept.

hero (hero):

But there is a y-intercept

OpenStudy (anonymous):

So the x intercept is 0? whats the y intercept?

hero (hero):

You are confused. I never said what any intercept is.

hero (hero):

What I posted were the steps to find the intercepts.

hero (hero):

y-intercept means that we automatically assume x will be zero because in order for a point on the graph to intercept the y-axis, x must be zero. To figure out at what value the point on the graph intercepts the y-axis, we input x = 0 into the given formula like so: \[y = \frac{2}{0^2 - 2(0) - 3}\]

hero (hero):

Notice that upon doing so, \[y = -\frac{2}{3}\]

hero (hero):

Where are you getting -2/5

hero (hero):

Do you have a 5 instead of 3 in your original problem?

OpenStudy (anonymous):

No, i realized what i was doing wrong. So the y intercept is -2/-3?

hero (hero):

I'm shaking my head right now. You've added an extra negative for no reason.

hero (hero):

The fraction is negative once. you don't add two negatives to a negative fraction. If you do that, the fraction becomes positive.

hero (hero):

\[-\frac{2}{3} \ne \frac{-2}{-3}\]

OpenStudy (anonymous):

Sorry, thats what i means, not two negatives.

OpenStudy (anonymous):

So the y intercept is (0,-2/3)

hero (hero):

Yes

hero (hero):

Now, if you've programmed your calculator as I have, all you have to do is just simply plot the points: f(-7) f(-6) f(-5) f(-4) f(-3) f(-2) f(0) f(1) f(2) f(4) f(5) f(6) f(7)

OpenStudy (anonymous):

Is the x intercept 0 or none? and i also need the domain

hero (hero):

To find the x intercept, let y = 0 then solve for x. If you are not able to solve for x, then there is no x-intercept.

hero (hero):

When you are done, you should get something like this: